# Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation

### v1

@@ Abstract
This study considers the estimation of the average treatment effect (ATE). For ATE estimation, we estimate the propensity score through direct bias-correction term estimation. Let $\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ be the observations, where $X_{i}\in{\mathbb{R}}^{K}$ denotes $K$-dimensional covariates, $D_{i}\in\{0,1\}$ denotes a binary treatment assignment indicator, and $Y_{i}\in{\mathbb{R}}$ is an outcome. In ATE estimation, the bias-correction term $h_{0}(X_{i},D_{i})\coloneqq\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}$ plays an important role, where $e_{0}(X_{i})$ is the propensity score, the probability of being assigned treatment $1$. In this study, we propose estimating $h_{0}$ (or equivalently $e_{0}$) by directly minimizing the prediction error for $h_{0}$ without knowing $h_{0}$ in advance. After showing a basic result with least squares, we present a general framework for this direct bias-correction term estimation approach from the perspective of Bregman divergence minimization, which also generalizes the Riesz regression and covariate balancing.
@@ 1 Introduction
In ATE estimation, the propensity score, the probability of treatment assignment given covariates, plays a crucial role (Rosenbaum & Rubin, 1983). For example, the inverse probability weighting (IPW) estimator, also known as the Horvitz-Thompson estimator (Horvitz & Thompson, 1952), estimates the ATE by computing a weighted average of outcomes using the inverse of the propensity score. In the one-step bias correction method (van der Vaart, 2002), a bias-correction term depending on the propensity score is added to an initial estimator.
In this study, we focus on the role of the propensity score as a bias-correction component and propose a novel method to directly estimate either the propensity score or the bias-correction term itself. The bias-correction term is fundamental to IPW-type estimators and the one-step bias correction method, and its accurate estimation can substantially improve ATE estimation. Importantly, our objective is not to estimate the propensity score, but rather the bias-correction term, in which the propensity score appears inversely and is weighted by the treatment indicator. Since estimating the propensity score itself is not the target, we hypothesize that directly estimating the bias-correction term will lead to improved performance. To this end, we propose estimating the bias-correction term by minimizing the empirical risk targeted for the true bias-correction term.
The technical challenge is that the target variable remains unobserved, even when we directly estimate the bias-correction term or the propensity score. To address this issue, we employ techniques developed in the direct density-ratio estimation (DRE) literature (Sugiyama et al., 2012). In direct DRE, the goal is to minimize the empirical risk between the true density ratio and its model, even though the true density ratio is unknown. It is known that empirical risk minimization is feasible even without knowledge of the true propensity score. Since the inverse propensity score can be viewed as a density ratio, we can extend these existing methods to our setting.
Our motivation is closely aligned with studies on covariate balancing (Imai & Ratkovic, 2013) and Riesz regression (Chernozhukov et al., 2022a), which also aim to improve ATE estimation by appropriately estimating the propensity score or the bias-correction term. Studies in covariate balancing focus on the balancing property of propensity score estimator and estimate them using the property. Chernozhukov et al. (2022a) proposes Riesz regression which represents the bias-correction term as the Riesz representer. Although the derivation process is different, we derive the objective function that is the same as Chernozhukov et al. (2022a) by using the DRE techniques. Further, we generalize our objective by using the Bregman divergence as well as DRE in Sugiyama et al. (2011). From this generalization, we connect our approach to the covariate balancing by showing the equivalence between our objective and empirical balancing (Chan et al., 2015).
@@ 1.1 ATE estimators and bias correction
Let $e_{0}(X)=P_{0}(D=1\mid X)$ denote the probability of assigning treatment $1$ given covariates $X$, which is known as the propensity score. Throughout this study, we impose the following condition, commonly referred to as the overlap assumption:
@@
There exists a constant $C>0$ independent of $n$ such that $C<e_{0}(x)<1-C$ for all $x\in\mathcal{X}$.
@@
$h_{0}(D,X)\coloneqq h(D,X;1/e_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}$
is crucial. This term, referred to as the bias-correction term, is central to ATE estimation (Schuler & van der Laan, 2024). A common approach is to estimate $e_{0}$ using logistic regression and then plug the resulting estimate $\widehat{e}^{{\mathrm{L}}}_{n}$ into $h$. Note that in automatic debiased machine learning, the term is also referred to as the Riesz representer (Chernozhukov et al., 2022b).
For example, in a typical one-step bias correction, we first construct an ATE estimator as $\widehat{\tau}^{\mathrm{DM}}_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\mu}_{n}(1,X)-\widehat{\mu}_{n}(0,X)\right)$, where $\widehat{\mu}_{n}$ is an estimator of $\mu_{0}$. This estimator is known as the direct method (DM) or naive plug-in estimator. To obtain an efficient estimator, we add the bias-correction term $\frac{1}{n}\sum^{n}_{i=1}h_{0}(X_{i},D_{i})$ to the first-stage DM estimator $\widehat{\tau}^{\mathrm{DM}}_{n}$, yielding the AIPW estimator.
In this process, estimating the propensity score $e_{0}$ itself is a challenging problem, independent of ATE estimation. The well-known Vapnik principle states that “when solving a problem of interest, do not solve a more general problem as an intermediate step” (Vapnik, 1998). Following this principle, this study aims to estimate $h_{0}(D,X)$, or equivalently $e_{0}$, by directly minimizing the estimation error of $h_{0}(D,X)$. We emphasize that while $e_{0}$ is estimated in our approach, the primary estimation target is not $e_{0}$ but $h_{0}$.
@@ 1.2 Our contributions
Our first contribution is the proposal of a framework for direct bias-correction term estimation. We estimate the bias-correction term by directly minimizing the estimation error for the true bias-correction term $h_{0}$. To model the bias-correction term $h_{0}$, we define the inverse propensity score as $r_{0}(1,X)=\frac{1}{e_{0}(X)}$ and $r_{0}(0,X)=\frac{1}{1-e_{0}(X)}$. Let $\mathcal{R}$ be a model (hypothesis class) for $r_{0}$, and consider approximating $r_{0}$ using some $r\in\mathcal{R}$. We approximate $r_{0}$ by the minimizer $r^{*}\in\mathcal{R}$ of the mean squared error (MSE) ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$. Here, we estimate $h_{0}(D,X)=h(D,X;r_{0})$ using a model $h(D,X;r)$ that depends on $r$. We emphasize again that although we bypass the estimation of $r_{0}$, our primary estimation target is $h_{0}$, not $r_{0}$.
Since this expected squared error involves the unknown function $r_{0}$, direct optimization is infeasible. However, we show that minimizing this expected squared error with respect to $r$ is equivalent to minimizing ${\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$, which does not depend on the unknown function. That is, we establish the equivalence: $r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$. The obtained squared error can then be approximated using a sample-based empirical risk function.
@@
Our second main contribution is the theoretical analysis of the estimator obtained via direct bias-correction term estimation. Since we estimate $r_{0}$ using empirical risk minimization, we establish bounds on the estimation error using empirical process theory. Furthermore, we present examples of ATE estimators that incorporate the bias-correction term estimated using our framework and conduct simulation studies. Using standard ATE estimation techniques, we demonstrate that our method yields a $\sqrt{n}$-consistent ATE estimator.
Our third main contribution is the generalization of our framework. From the perspective of Bregman divergence minimization, we extend our framework to provide a more general methodology for direct bias-correction term estimation. Under this framework, various estimation strategies can be incorporated to enhance bias-correction term estimation.
Our fourth contribution is the unification of the existing literature: Riesz regression and covariate balancing. If we use the squared loss function in the Bregman divergence minimization, we obtain the same loss as the one used in Riesz regression. If we use a different function in the Bregman divergence, we recover a special case of the tailored loss for covariate balancing proposed in Zhao (2019), which is also equivalent to empirical balancing (Chan et al., 2015; Hainmueller, 2012).
2 Direct bias-correction term estimation
In this study, we consider estimating $h_{0}$ by minimizing the empirical risk associated with the MSE ${\mathbb{E}}\left[\big(h_{0}(D,X)-h(D,X)\big)^{2}\right]$ over $h:\{1,0\}\times\mathcal{X}\to{\mathbb{R}}$. Here, since $h_{0}$ contains the unknown function $r_{0}$, this minimization problem is inherently intractable. However, we demonstrate that empirical risk minimization remains feasible without explicit knowledge of $r_{0}$.
2.1 Least squares
Rather than directly modeling $h_{0}$, we model either $r_{0}$ or $e_{0}$ within $h_{0}$. That is, we model $h_{0}$ via a model of $r_{0}$ as $h(D,X;r)$, while $h_{0}(D,X)$ is given as $h_{0}(D,X;r_{0})$. We then aim to estimate $r_{0}$ or $e_{0}$. Given a set $\mathcal{R}$ of functions $r\colon{\mathcal{X}}\to(1,\infty)$, we approximate the bias-correction term by solving
$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right].$
If $r_{0}\in\mathcal{R}$, then it follows that $r_{0}(\cdot,\cdot)=r^{*}(\cdot)$ and $h(\cdot,\cdot;r^{*})=h_{0}(\cdot,\cdot)$. This MSE is associated with the MSE between the ATE $\tau_{0}$ and some estimator using the bias-correction term. Since $e_{0}(x)\in(0,1)$, a model $r(x)$ takes a value in $(1,\infty)$.
Example. For example, for $\widehat{\tau}^{\mathrm{IPW}}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}h(D_{i},X_{i};r)Y_{i}$, we have ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}+\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}\right)^{2}\right]+{\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$, where recall that we defined $\widetilde{\tau}^{\mathrm{IPW}}$ in the Introduction. Here, we have ${\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]=\frac{1}{n}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}{\mathbb{E}}\left[Y^{2}\mid D,X\right]\right]$. Thus, the MSE ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$ is closely connected to the MSE ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$.
Direct estimation without $r_{0}$. The essential problem in this approach is that the target $h_{0}$ or $r_{0}$ is unknown; therefore, this optimization looks infeasible. However, even if we do not know $h_{0}$, we can reformulate the optimization problem into an equivalent form that does not involve $h_{0}$ as
$r^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\Big[-2r(1,X)-2r(0,X)+\mathbbm{1}[D=1]r(1,X)^{2}+\mathbbm{1}[D=0]r(0,X)^{2}\Big].$
This reformulation is derived as follows: $\min_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\Big[h(D,X;r_{0})^{2}-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\Big]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$. From the second to the third line, we omit the term $h_{0}(D,X)^{2}$ since it does not affect the optimization. From the third to the fourth line, we use the following for the first term: ${\mathbb{E}}\left[h(D,X;r_{0})h(D,X)\right]={\mathbb{E}}\Bigg[\left(\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}\right)\Big(\mathbbm{1}[D=1]r(1,X)-\mathbbm{1}[D=0]r(0,X)\Big)\Bigg]={\mathbb{E}}\left[\frac{\mathbbm{1}[D=1]}{e_{0}(X)}r(1,X)+\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[\frac{e_{0}(X)}{e_{0}(X)}r(1,X)+\frac{1-e_{0}(X)}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[r(1,X)+r(0,X)\right]$.

### v2

# Direct Bias-Correction Term Estimation for Average Treatment Effect Estimation

@@ Abstract
This study considers the estimation of the direct bias-correction term for estimating the average treatment effect (ATE). Let $\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ be the observations, where $X_{i}\in{\mathbb{R}}^{K}$ denotes $K$-dimensional covariates, $D_{i}\in\{0,1\}$ denotes a binary treatment assignment indicator, and $Y_{i}\in{\mathbb{R}}$ denotes an outcome. In ATE estimation, $h_{0}(D_{i},X_{i})\coloneqq\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}$ is called the bias-correction term, where $e_{0}(X_{i})$ is the propensity score. The bias-correction term is also referred to as the Riesz representer or clever covariates, depending on the literature, and plays an important role in construction of efficient ATE estimators. In this study, we propose estimating $h_{0}$ by directly minimizing the Bregman divergence between its model and $h_{0}$, which includes squared error and Kullback–Leibler divergence as special cases. Our proposed method is inspired by direct density ratio estimation methods and generalizes existing bias-correction term estimation methods, such as covariate balancing weights, Riesz regression, and nearest neighbor matching. Importantly, under specific choices of bias-correction term models and Bregman divergence, we can automatically ensure the covariate balancing property. Thus, our study provides a practical modeling and estimation approach through a generalization of existing methods.
@@ 1.1 ATE estimators and bias correction
Let $e_{0}(X)=P_{0}(D=1\mid X)$ denote the probability of assigning treatment $1$ given covariates $X$, which is known as the propensity score. Throughout this study, we impose the following conditions, commonly referred to as the unconfoundedness and common support assumptions.
@@
[…]
@@
$h_{0}(D,X)\coloneqq h(D,X)\coloneqq\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}$
is crucial. This term, referred to as the bias-correction term, is central to ATE estimation (Schuler & van der Laan, 2024). A common approach is to estimate $e_{0}$ using logistic regression and then plug the resulting estimate $\widehat{e}^{{\mathrm{L}}}_{n}$ into $h$. Note that the bias-correction term is also referred to as the Riesz representer (Chernozhukov et al., 2021) or the clever covariates (van der Laan, 2006). We use the term bias-correction term because the Riesz representer is closely connected to the automatic debiased machine learning literature, and the clever covariates is closely connected to the targeted maximum likelihood estimation (TMLE) literature.
For example, in a typical one-step bias correction, we first construct an ATE estimator as $\widehat{\tau}^{\mathrm{DM}}_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\mu}_{n}(1,X)-\widehat{\mu}_{n}(0,X)\right)$, where $\widehat{\mu}_{n}$ is an estimator of $\mu_{0}$. This estimator is known as the direct method (DM) or naive plug-in estimator. To obtain an efficient estimator, we add the bias-correction term $\frac{1}{n}\sum^{n}_{i=1}h_{0}(D_{i},X_{i})(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i}))$ to the first-stage DM estimator $\widehat{\tau}^{\mathrm{DM}}_{n}$, yielding the AIPW estimator.
In this study, we propose a method to estimate the bias-correction term, also called the Riesz representer or the clever covariates. For example, we can estimate the bias-correction term by estimating the propensity score $e_{0}$ using the maximum likelihood estimation. However, our interest is not in propensity score estimation but in bias-correction term estimation. As the well-known Vapnik principle states, we should avoid such an intermediate problem and ideally aim to estimate the target objective in a more direct manner (Vapnik, 1998). Following this principle, this study considers estimating $h_{0}(D,X)$ by directly minimizing the estimation error for the true $h_{0}(D,X)$.
The technical challenge is that the target objective $h_{0}$ is unknown. To address this issue, we employ techniques developed in the direct density-ratio estimation (DRE) literature (Sugiyama et al., 2012). In direct DRE, the goal is to minimize the empirical risk between the true density ratio and its model, even though the true density ratio is unknown. It is known that empirical risk minimization is feasible even without knowledge of the true propensity score. Since the inverse propensity score can be viewed as a density ratio, we can extend these existing methods to our setting. For causal inference researchers who are unfamiliar with DRE, we review the DRE literature in Appendix A.
Our motivation is also closely aligned with studies on Riesz regression (Chernozhukov et al., 2021) and covariate balancing weights (Imai & Ratkovic, 2013b; Deville & Särndal, 1992), which also aim to estimate the bias-correction term in a direct manner. Studies in covariate balancing focus on the balancing property of propensity score estimator and estimate them using the property. Chernozhukov et al. (2021) proposes Riesz regression which represents the bias-correction term as the Riesz representer. Although the derivation process is different, we derive the objective function that is the same as Chernozhukov et al. (2021) by using the DRE techniques. Further, we generalize our objective by using the Bregman divergence as well as DRE in Sugiyama et al. (2011). From this generalization, we further connect our approach to the covariate balancing by showing the equivalence between our objective and empirical balancing through the duality arguments discussed in Zhao (2019) and Bruns-Smith et al. (2025).
@@ 1.2 Our contributions
This study has the following four contributions: (i) a general framework for directly estimating the bias-correction term (also called the Riesz representer or clever covariates) via Bregman divergence minimization; (ii) our proposed framework includes Riesz regression in Chernozhukov et al. (2021) and the tailored loss in Zhao (2019) as special cases; (iii) under our framework, we show that there are appropriate choices of bias-correction term models and Bregman divergences under which covariate balancing is automatically realized as the dual of the Bregman divergence minimization problem (automatic covariate balancing); (iv) we provide a theoretical analysis of the estimator.
Our first contribution is the proposal of a framework for direct bias-correction term estimation via Bregman divergence minimization. We estimate the bias-correction term by directly minimizing the estimation error of the true bias-correction function $h_{0}$, measured by the Bregman divergence, $\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)\coloneqq{\mathbb{E}}\Big[g\big(h_{0}(D,X)\big)-g\big(h(D,X)\big)-\partial g\big(h(d,X)\big)\big(h_{0}(D,X)-h(D,X)\big)\Big]$, where $g$ is a differentiable and strictly convex function. By changing $g$, we can measure the error using various metrics, such as the squared loss or KL divergence loss. Since the Bregman divergence involves the unknown function $h_{0}$, direct optimization is infeasible. To address this issue, we propose minimizing an alternative objective function, defined as $\mathrm{BR}_{g}\big(h\big)\coloneqq{\mathbb{E}}\Big[-g\big(h(D,X)\big)+\partial g\big(h(D,X)\big)h(D,X)-\partial g\big(h(1,X)\big)-\partial g\big(h(0,X)\big)\Big]$. Minimizing the original Bregman divergence $\mathrm{BR}^{\dagger}_{g}\big(h\big)$ is equivalent to minimizing $\mathrm{BR}_{g}\big(h\big)$, which does not depend on the unknown function. That is, we establish the equivalence: $h^{*}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g}\big(h\big)$. The resulting objective function can then be approximated using an empirical risk function.
Our second contribution is the unification of existing literature. Our proposed Bregman divergence minimization objective includes Riesz regression in Chernozhukov et al. (2021) (when using the squared loss) and the tailored loss in Zhao (2019) (when using the KL divergence loss). Furthermore, our framework also integrates covariate balancing methods (Imai & Ratkovic, 2013a; Hainmueller, 2012; Zubizarreta, 2015; Chan et al., 2015; Wong & Chan, 2017). If we use linear models to approximate the bias-correction term and train the model with the squared loss (Riesz regression), the dual problem coincides with the optimization problem in stable balancing weights. If we model the bias-correction term via the propensity score with logistic models and train the model with the KL divergence loss (tailored loss), the dual problem becomes the same as the optimization problem in entropy balancing weights. Kato (2025a), a subsequent work of this study, refers to this property as automatic covariate balancing. See Table 1 in Section 2 and Figure 1 in Appendix.
@@
Our third main contribution is the theoretical analysis of the estimator obtained via direct bias-correction term estimation. Since we estimate $r_{0}$ using empirical risk minimization, we establish bounds on the estimation error using empirical process theory. Furthermore, we present examples of ATE estimators that incorporate the bias-correction term estimated using our framework and conduct simulation studies. Using standard ATE estimation techniques, we demonstrate that our method yields a $\sqrt{n}$-consistent ATE estimator.
As a side product of our contributions, we find that we can import various existing results from the DRE literature. Since Riesz regression is essentially the same as LSIF, various results about convergence rate analysis and optimization methods have already been established. For example, Kanamori et al. (2012) shows the convergence rate when using a reproducing kernel hilbert space (RKHS) for the density ratio, or equivalently the bias-correction term. Kato & Teshima (2021) shows the rate when using neural networks, which has been further refined in Zheng et al. (2022). Rhodes et al. (2020) and Kato & Teshima (2021) point out the overfitting problem characteristic of DRE estimation and propose techniques to avoid the problem. Lin et al. (2023) finds that nearest neighbor matching can be interpreted as density ratio estimation, and it can also be interpreted as a special case of LSIF or Riesz regression (See Appendix H). These findings not only help deepen our understanding of Riesz regression, but also prevent unnecessary reinvention. For example, the covariate adaption method proposed in Chernozhukov et al. (2025) uses Riesz regression, but it is essentially the same as covariate adaption with a density ratio estimated via LSIF (Kanamori et al., 2009), except for the regression adjustment. While Chernozhukov et al. (2022a) proposes neural networks and random forests for Riesz regression, the techniques for estimating the density ratio have also been proposed in the DRE literature (Kanamori et al., 2012; Abe & Sugiyama, 2019; Rhodes et al., 2020; Kato & Teshima, 2021).
2 Bias-correction term estimation via Bregman divergence minimization
In this study, we consider estimating $h_{0}$ by minimizing the empirical risk associated with the Bregman divergence between $h_{0}$ and its estimator $h\colon\{1,0\}\times\mathcal{X}\to{\mathbb{R}}$.
2.1 Population Bregman divergence minimization
Let $g\colon{\mathbb{R}}\to{\mathbb{R}}$ be a differentiable and strictly convex function. Given $d\in\{1,0\}$, we define the Bregman divergence between $h_{0}$ and $h$ as $\mathrm{br}^{\dagger}_{g}\big(h_{0}(d,x)\mid h(d,x)\big)\coloneqq g\big(h_{0}(d,x)\big)-g\big(h(d,x)\big)-\partial g\big(h(d,x)\big)\big(h_{0}(d,x)-h(d,x)\big)$, where $\partial g$ denotes the derivative of $g$. Then, we define the average Bregman divergence as $\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)\coloneqq{\mathbb{E}}\Big[g\big(h_{0}(D,X)\big)-g\big(h(D,X)\big)-\partial g\big(h(d,X)\big)\big(h_{0}(D,X)-h(D,X)\big)\Big]$. Then, we estimate $h_{0}$ by $h^{*}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)$. By dropping the term that is irrelevant to learning, we have
$h^{*}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g}\big(h\big),$
$\text{where}\quad\mathrm{BR}_{g}\big(h\big)\coloneqq{\mathbb{E}}\Big[-g\big(h(D,X)\big)+\partial g\big(h(D,X)\big)h(D,X)-\partial g\big(h(1,X)\big)+\partial g\big(h(0,X)\big)\Big].$
This can be shown as follows:
$h^{*}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\left[\mathbbm{1}[D=d]\left(g(h_{0}(d,X))-g(h(d,X))-\partial g(h(d,X))\Big(h_{0}(d,X)-h(d,X)\Big)\right)\right]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{H}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(h(d,X))-\partial g(h(d,x))\Big(h_{0}(d,X)-h(d,X)\Big)\right)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{H}}}\sum_{d\in\{1,0\}}\left({\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(h(d,X))+\partial g(h(d,X))h(d,X)\right)\Big]-{\mathbb{E}}\Big[\mathbbm{1}[D=d]\partial g(h(d,x))h_{0}(d,X)\Big]\right)$
$=\operatorname*{arg\,min}_{r\in{\mathcal{H}}}\left\{{\mathbb{E}}\Big[\left(-g(h(D,X))+\partial g(h(D,X))h(d,X)\right)\Big]-{\mathbb{E}}\Big[\partial g(h(1,X))\Big]+{\mathbb{E}}\Big[\partial g(h(0,X))\Big]\right\}.$
Here, we dropped terms irrelevant to the optimization and used ${\mathbb{E}}[\mathbbm{1}[D=1]h_{0}(1,X)\mid X]={\mathbb{E}}[e_{0}(X)h_{0}(1,X)\mid X]=1$ and ${\mathbb{E}}[\mathbbm{1}[D=0]h_{0}(0,X)\mid X]=-1$.